# More Experts Than Galaxies: Conditionally-Overlapping Experts with Biologically-Inspired Fixed Routing

### v2

@@ 4.2.1 STANDARD MLP - CIFAR10
MoE Trainable: A MoE model with a trainable routing function, where the number of experts is equal to $\lfloor 1/p_{\rm k}\rfloor$, and each expert has $p_{\rm k}N_{\ell}$ neurons in each layer. The routing network is a trainable MLP with one hidden layer, whose output is a sparse $\lfloor 1/p_{\rm k}\rfloor$-dimensional vector.
@@
Bernoulli Masking: An MLP where each example in the training data is associated with a fixed binary mask drawn from a Bernoulli distribution, with probability equal to $p_{\rm k}$. Thus the relationship between inputs and their masks is arbitrary, rather than being mediated by the routing network in COMET.
Example-tied Dropout: Example-tied dropout (Maini et al., 2023), where each example in the training data is associated with a fixed binary mask drawn from a Bernoulli distribution, with probability equal to $p_{\rm k}$, and a fixed number of “generalization neurons” are active for all examples (reduces to Bernoulli Masking when there are zero generalization neurons).
@@
We evaluate these models on the CIFAR10 dataset Krizhevsky (2009), with results shown in Figure 4. Overall, the optimal model architecture depends on the capacity of the network. When the number of neurons is limited and the network has a low capacity to learn the task (i.e., low $p_{\rm k}$), the standard model that utilizes all neurons outperforms most models. However, as network capacity increases with more neurons, the COMET model emerges as the top performer. This suggests that the benefits of selective neuron activation become more pronounced as capacity increases.
Figure 4: Illustration of 4-layer MLP networks trained on CIFAR10, showcasing the impact of varying network capacity and sparsity levels. As we systematically increase the number of neurons and decrease sparsity (moving from top left to bottom right), we observe a shift in the best-performing model. Initially, the standard model outperforms the COMET model when network capacity is low. However, as network capacity grows, the COMET model emerges as the top performer.
@@ 4.2.2 CONTEMPORARY ARCHITECTURES
We further extend the COMET method to contemporary architectures in the Vision domain, including ViT Dosovitskiy et al. (2021) and MLP-Mixer Tolstikhin et al. (2021). To do this, we apply the COMET random projection followed by the cap operation in the MLP layers of each with $p_{\rm k}=0.5$. We evaluate the performance of these models on four widely-used image classification datasets: SVHN Netzer et al. (2011), CIFAR10 Krizhevsky (2009), CIFAR100 Krizhevsky (2009), and Tiny ImageNet Le & Yang (2015). Our results can be seen in Figures 6, 6, 8, 8.
@@
4.3 LANGUAGE MODELING
We apply COMET to language modeling on Wikitext Merity et al. (2016) and CodeParrot Tunstall et al. (2022) with varying GPT model sizes, with results in Figures 10 and 10. We again observe the same pattern: as network capacity increases, the COMET model outperforms the standard model, with larger models exhibiting not only a more significant performance difference but also faster learning rates, highlighting the benefits of selective neuron activation in language modeling tasks.
4.4 REGRESSION
To further validate our results, we also evaluated COMET on the SARCOS regression dataset. Our findings in Appendix A.5 show that our conclusions generalize to this setting as well.
@@ 5 CONCLUSIONS
In this work, we propose a sparse neural network method, COMET, that induces a modular, sparse architecture with an exponential number of overlapping experts and alleviates key limitations of existing modular approaches. Specifically, we tackle the issues of trainable gating functions that often lead to representation collapse, non-overlapping experts that hinder knowledge transfer, and the need for explicit input IDs that can be restrictive. By leveraging a fixed, biologically-inspired random projection, COMET determines expert overlap based on input similarity in a fully unsupervised manner, enabling faster learning and improved generalization through enhanced forward transfer. Through extensive experiments on various tasks, including image classification, language modeling, and regression, we demonstrate that COMET achieves improved performance, especially for larger models, demonstrating that our method is applicable to many popular model architectures.
@@ Appendix A EXPERIMENT DETAILS
A.1 MAIN RESULTS: STANDARD MLP - CIFAR10
@@
A.2 ADDITIONAL RESULTS: STANDARD MLP - CIFAR10
@@
As illustrated in Figure 11, a consistent trend emerges as we systematically vary the number of neurons and sparsity levels. Moving from top left to bottom right, we observe a shift in the optimal model configuration. Initially, when network capacity is limited, the standard model outperforms the COMET model. However, as network capacity increases, the COMET becomes the top performer, surpassing every other model. This trend reinforces our key finding: selective neuron activation becomes increasingly beneficial as network capacity increases, enabling faster learning and improved generalization through enhanced forward transfer.
@@ A.3 CONTEMPORARY ARCHITECTURES
We built upon the tuned hyperparameters from Yoshioka (2024) and made the following modifications to analyze the effect of our sparsity method, COMET. Specifically, we systematically increased model capacity by adding more neurons to the MLP layers of each model. Additionally, we used the $\tanh$ activation function on the backbone MLP layers. See Appendix A.6 for further analysis on activation functions.
@@
A.4 LANGUAGE MODELING
@@ A.4.2 Additional Results
To validate our main findings, we first assess the performance of the different GPT-2 model sizes on WikiText and CodeParrot using the standard GPT-2 tokenizer and $50\%$ sparsity level. Our results are presented in Figures 12, 13, 14, 15, 16, and 17.
@@
To investigate the impact of learning rate on the COMET method, we scale the learning rate by a factor of 10, from 1e-4 to 1e-3. The results of this experiment are presented in Figures 18, 19, 20, 21, 22, and 23.
@@
Furthermore, we identify an interesting trend as the model size increases, as seen in Figures 19, 20, 22, and 23. Specifically, when using a larger learning rate and larger model sizes, the standard models tend to overfit or experience exploding gradients, resulting in a significant increase in validation perplexity. In contrast, adding sparsity using COMET not only enables faster learning but also mitigates overfitting and gradient explosion, allowing for stable training with a larger learning rate.
@@
To further investigate the robustness of the COMET method, we examine the effect of reducing the batch size by a factor of 4, achieved by decreasing the gradient accumulation step from 8 to 2. Our results are presented in Figures 24, 25, 26, 27, 28, and 29.
@@
We further investigate the robustness of the COMET method by evaluating its performance under mixed precision training. Specifically, we assess the impact of switching from FP16 to FP32 precision on the COMET-based models. Our results are presented in Figures 30, 32, and 31.
@@ Results
The results, presented in Figure 33, demonstrate the effectiveness of our approach in this domain. Consistent with our previous findings, we again observe that the COMET model outperforms the standard model as network capacity increases, confirming the benefits of selective neuron activation in regression tasks.
A.6 ACTIVATIONS ANALYSIS
@@
Our results are presented in Figure 34. We find that the COMET model outperforms the standard model when most activation functions are applied on the backbone network. However, we observe that COMET does not perform as well as the standard model for certain non-monotonic activation functions, specifically GELU, Mish, and SiLU, under the settings we chose (sparsity level = 0.5 and 3000 neurons in each layer). Due to time constraints, we do not delve deeper into the reasons behind this phenomenon but suggest that future work could investigate why non-monotonic functions may hinder the effectiveness of COMET.
@@
A.7 COMET: Routing Network Architecture

### v4

@@ 5.2.1 Standard MLP – CIFAR10
MoE Trainable: A MoE model with $\lfloor 1/p_{\rm k}\rfloor$ experts, each having $p_{\rm k}N_{\ell}$ neurons in each layer. The routing network is a trainable MLP with one hidden layer and a sparse $\lfloor 1/p_{\rm k}\rfloor$-dimensional output.
@@
Bernoulli Masking: An MLP where each training example is associated with a fixed binary mask drawn from a Bernoulli distribution, with probability equal to $p_{\rm k}$. Thus the relationship between inputs and their masks is arbitrary, rather than being mediated by the routing network in COMET.
Example-tied Dropout: Example-tied dropout (Maini et al., 2023), where each example in the training data is associated with a fixed binary mask drawn from a Bernoulli distribution, with probability equal to $p_{\rm k}$, and a fixed number of “generalization neurons” are active for all examples.
@@
We evaluate these models on the CIFAR10 dataset Krizhevsky (2009), with results shown in fig. 4. Overall, the optimal model architecture depends on the capacity of the network. When the number of neurons is limited and the network has a low capacity to learn the task (i.e., low $p_{\rm k}$), the standard model that utilizes all neurons outperforms most models. However, as network capacity increases with more neurons, the COMET model emerges as the top performer. This suggests that the benefits of selective neuron activation become more pronounced as capacity increases. Notice in the high-capacity regime the Smaller Model matches the Standard Model, indicating that simply adding more neurons does not improve performance while adding neurons subject to COMET’s structured sparsity does.
Figure 4: Illustration of 4-layer MLP networks trained on CIFAR10, showcasing the impact of varying network capacity and sparsity levels. As we increase the number of neurons and decrease sparsity (moving from top left to bottom right), we observe a shift in the best-performing model. Initially, the standard model outperforms the COMET model when network capacity is low. However, as network capacity grows, the COMET model emerges as the top performer.
@@ 5.2.2 Contemporary Architectures
We further extend the COMET method to contemporary architectures in the Vision domain, including ViT Dosovitskiy et al. (2021) and MLP-Mixer Tolstikhin et al. (2021). To do this, we apply the COMET random projection followed by the cap operation in the MLP layers of each with $p_{\rm k}=0.5$. We evaluate the performance of these models on four widely-used image classification datasets: SVHN Netzer et al. (2011), CIFAR10 Krizhevsky (2009), CIFAR100 Krizhevsky (2009), and Tiny ImageNet Le & Yang (2015). Our results can be seen in figs. 11, 11, 6 and 6.
@@
5.3 Language Modeling and Regression
We apply COMET to language modeling on Wikitext (Merity et al., 2016) and CodeParrot (Tunstall et al., 2022) with varying GPT model sizes, with results in figs. 8 and 8. We again observe that as network capacity increases, the COMET model outperforms the standard model, with larger models exhibiting not only a greater performance difference but also faster learning rates, highlighting the benefits of selective neuron activation in language modeling tasks. To further validate our results, we also evaluated COMET on the SARCOS regression dataset. Our findings in section A.6 show that our conclusions generalize to this setting as well.
@@ 6 Conclusions
In this work, we propose a sparse neural network method, COMET, that induces a modular, sparse architecture with an exponential number of overlapping experts and alleviates key limitations of existing modular approaches, including trainable gating functions that often lead to representation collapse, non-overlapping experts that hinder knowledge transfer, and the need for explicit input IDs. By leveraging a biologically-inspired fixed random projection and $k$-winner-take-all capping operation, COMET determines expert overlap based on input similarity in a fully unsupervised manner, enabling faster learning and improved generalization through enhanced forward transfer. It should be noted that key features of COMET may occur naturally in standard networks. First, similar inputs already tend to have similar hidden representations (see the $p_{\rm k}=1$ curve in fig. 2). Second, there may be cases where networks self-organize to process different inputs in different subspaces of activation space (see elhage2022toy, for a related demonstration). Nevertheless COMET magnifies these effects and our experiments show it improves performance over standard networks. Through extensive experiments on various tasks, including image classification, language modeling, and regression, we demonstrate that COMET achieves improved performance, especially for larger models, demonstrating that our method is applicable to many popular model architectures.
Acknowledgements
This research was supported by the Intramural Research Program of the National Institute of Mental Health (NIMH), under project ZIC-MH002968 (PI: Dr. Francisco Pereira), and grant 2020906 from the National Science Foundation (PI: Matt Jones). We would also like to thank Michael Mozer for his invaluable insights and guidance.
@@ Appendix A Experiment Details
A.1 Main Results: Standard MLP – CIFAR10
@@
A.2 Additional Results: Standard MLP – CIFAR10
@@
As illustrated in fig. 9, a consistent trend emerges as we systematically vary the number of neurons and sparsity levels. Moving from top left to bottom right, we observe a shift in the optimal model configuration. Initially, when network capacity is limited, the standard model outperforms the COMET model. However, as network capacity increases, the COMET becomes the top performer, surpassing every other model. This trend reinforces our key finding: selective neuron activation becomes increasingly beneficial as network capacity increases, enabling faster learning and improved generalization through enhanced forward transfer.
@@ A.3 Contemporary Architectures
We built upon the tuned hyperparameters from Yoshioka (2024) and made the following modifications to analyze the effect of our sparsity method, COMET. Specifically, we systematically increased model capacity by adding more neurons to the MLP layers of each model. Additionally, we used the $\tanh$ activation function on the backbone MLP layers. See section A.7 for further analysis on activation functions.
@@
A.4 Image Classification
We provide additional experimental results on the image classification task, evaluated on the CIFAR-10 and SVHN datasets, as illustrated in Figure 11 and Figure 11, respectively.
A.5 Language Modeling
@@ A.5.2 Additional Results
To validate our main findings, we first assess the performance of the different GPT-2 model sizes on WikiText and CodeParrot using the standard GPT-2 tokenizer and $50\%$ sparsity level. Our results are presented in figs. 12, 13, 14, 15, 16 and 17.
@@
To investigate the impact of learning rate on the COMET method, we scale the learning rate by a factor of 10, from 1e-4 to 1e-3. The results of this experiment are presented in figs. 18, 19, 20, 21, 22 and 23.
@@
Furthermore, we identify an interesting trend as the model size increases, as seen in figs. 19, 20, 22 and 23. Specifically, when using a larger learning rate and larger model sizes, the standard models tend to overfit or experience exploding gradients, resulting in a significant increase in validation perplexity. In contrast, adding sparsity using COMET not only enables faster learning but also mitigates overfitting and gradient explosion, allowing for stable training with a larger learning rate.
@@
To further investigate the robustness of the COMET method, we examine the effect of reducing the batch size by a factor of 4, achieved by decreasing the gradient accumulation step from 8 to 2. Our results are presented in figs. 24, 25, 26, 27, 28 and 29.
@@
We further investigate the robustness of the COMET method by evaluating its performance under mixed precision training. Specifically, we assess the impact of switching from FP16 to FP32 precision on the COMET-based models. Our results are presented in figs. 30, 32 and 31.
@@ Results
The results, presented in fig. 33, demonstrate the effectiveness of our approach in this domain. Consistent with our previous findings, we again observe that the COMET model outperforms the standard model as network capacity increases, confirming the benefits of selective neuron activation in regression tasks.
A.7 Activations Analysis
@@
Our results are presented in fig. 34. We find that the COMET model outperforms the standard model when most activation functions are applied on the backbone network. However, we observe that COMET does not perform as well as the standard model for certain non-monotonic activation functions, specifically GELU, Mish, and SiLU, under the settings we chose (sparsity level = 0.5 and 3000 neurons in each layer). Due to time constraints, we do not delve deeper into the reasons behind this phenomenon but suggest that future work could investigate why non-monotonic functions may hinder the effectiveness of COMET.
@@
A.8 COMET: Routing Network Architecture
@@
A.9 Running Time and Memory Requirements
While COMET does not introduce additional trainable parameters, it does require some extra computation to calculate the masks. We evaluate the training times and GPU memory usage of COMET compared to standard vision models.
Tables 1 and 2 show that while COMET introduces some additional computation, the overhead is moderate. For example, the added GPU usage ranges from $0.3$ GB ($4\%$) to $7$ GB ($54\%$), where the larger percentage differences occur in cases where the ViT network is excessively overparameterized (e.g., when the number of neurons in the original network is multiplied by 12, from 512 to 6144).
It is worth noting that for simplification, our implementation has not been optimized:
  1. 1.
    We did not use sparse matrix multiplication methods, which could avoid unnecessary computations through the zero masks. This would reduce both the training time (by preventing gradient calculation of the zero gates) and overall GPU usage and inference time (due to less computation). This optimization could lead to significant improvements, potentially even compared to the base models.
  2. 2.
    As mentioned in section A.8, we focus on a specific case where the fixed matrix ${\bm{C}}_{i}$ has the same dimensions as the weight matrix ${\bm{W}}_{i}$ at each layer. This means that each random projection matrix has the same shape as the corresponding FC layer’s matrix. Notably, however, the only requirement is that they share the same input and output dimensions, leaving room for future exploration of alternative configurations.
Overall, these results suggest that COMET’s additional computational overhead is manageable, even with our simple implementation, and that its benefits can be achieved without significantly sacrificing efficiency.
Dataset / Model
SVHN CIFAR10 CIFAR100 Tiny ImageNet
COMET ViT 6.6 6.6 6.6 6.6
Standard ViT 6.1 6.1 6.1 6.1
COMET ViT-Med 12.3 12.3 12.3 12.3
Standard ViT-Med 9.2 9.2 9.2 9.2
COMET ViT-Large 19.8 19.8 19.8 19.8
Standard ViT-Large 12.8 12.8 12.8 12.8
COMET MLP-Mixer 6.7 6.7 6.7 6.7
Standard MLP-Mixer 6.4 6.4 6.4 6.4
COMET MLP-Mixer-Med 25.5 25.5 25.5 25.5
Standard MLP-Mixer-Med 23.2 23.2 23.2 23.2
Table 1: GPU Utilization in GBs. All models were trained on a single A100 GPU.
Dataset / Model
SVHN CIFAR10 CIFAR100 Tiny ImageNet
COMET ViT 1.1 0.8 0.8 1.2
Standard ViT 0.9 0.7 0.7 1.2
COMET ViT-Med 2.0 1.5 1.5 2.4
Standard ViT-Med 1.3 1.1 1.1 1.3
COMET ViT-Large 3.3 2.2 2.0 3.3
Standard ViT-Large 1.8 1.4 1.3 1.7
COMET MLP-Mixer 1.8 1.5 1.4 2.3
Standard MLP-Mixer 1.6 1.4 1.3 2.3
COMET MLP-Mixer-Med 10.2 5.4 6 9.3
Standard MLP-Mixer-Med 7.6 4.5 4.8 8.1
Table 2: Training times in hours. All models were trained on a single A100 GPU.
A.10 Transfer Learning
Figure 35: Non-IID Transfer: Evaluating COMET and the standard model on non-iid transfer learning. In particular, we train both COMET and the standard model on CIFAR10 followed by finetuning on SVHN (with transfer), or just train both models on SVHN (without transfer).
To further demonstrate COMET’s ability to transfer learned knowledge, we conducted an additional experiment. We trained MLP COMET and standard MLP models on CIFAR-10 and then fine-tuned them on SVHN. The results, presented in fig. 35, show that COMET exhibits superior transfer learning capabilities. Specifically, “COMET with transfer” outperforms “standard model with transfer” in both initial accuracy and learning speed. Moreover, “COMET with transfer” also surpasses the overall performance of both “COMET without transfer” and “standard model without transfer” on SVHN. This suggests that COMET’s advantage is not limited to out-of-sample generalization in the same distribution but also extends to transfer across tasks.
Interestingly, we observe a larger gap in performance between “standard model with transfer” and “standard model without transfer” compared to the gap between “COMET model with transfer” and “COMET model without transfer”. This might suggest that the standard model benefits more from transfer learning on non-iid data. However, since overall performance is generally taken as the primary metric for evaluating transfer learning, COMET’s superior performance on SVHN is the most relevant outcome. This finding is promising for future work that will more thoroughly investigate COMET in multi-task settings including transfer learning and continual learning.